OpenStudy (anonymous):

two complementary angles are in the ratio 3:6 find the measure of each angle

OpenStudy (anonymous):

3:6 reduces to 1:2 let x=lesser angle, then x+2x=90 3x=90 x=30 degrees the larger is twice the lesser or 60 degrees

myininaya (myininaya):

let x and y be the angles x/y=3/6 => x/y=1/2 =>x=y/2 x+y=90 y/2+y=90 y+2y=180 3y=180 => y=60 since y=60 then x=30 gj madolino

OpenStudy (anonymous):

the ratio of measure of two supplementary angels is 9:7 what the answer?